OpenStudy (anonymous):

Use the matrix method to solve the system of equations. A man has 30 coins, which are all dimes or nickels. Their combined value is $2.10. How many coins of each kind does he have? Let x = the number of nickels. Let y = the number of dimes. Set up two linear equations and use the matrix method to solve. Number of nickels =___ Number of dimes =___

OpenStudy (anonymous):

x + y = 30 0.05x + 0.1y = 2.1 Find the determinant of: \[\left[\begin{matrix}1 & 1 \\ 0.05& 0.1\end{matrix}\right]\] This is the system determinant. Find the determinant of: \[\left[\begin{matrix}30 & 1 \\ 2.1 & 0.1\end{matrix}\right]\] This is the x determinant. Find the determinant of: \[\left[\begin{matrix}1 & 30 \\ 0.05 & 2.1\end{matrix}\right]\] THis is the y determinant. Now divide the x determinant by the system determinant to get the x-value. Then divide the y determinant by the system determinant to get the y-value.

OpenStudy (anonymous):

I`m sorry but how do yo divide the system determinant?

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

Do you know how to find the system determinant?

OpenStudy (anonymous):

no :S

OpenStudy (anonymous):

1(0.1) - 1(0.05) = 0.05 You multiply diagonally and then subtract. It's diagonally down first, and then subtract by the upward diagonal.

OpenStudy (anonymous):

ad - bc

OpenStudy (anonymous):

Got it! so the number of nickels would be 18 and number of dimes would be dimes would be 12?

OpenStudy (anonymous):

Yupeee!

OpenStudy (anonymous):

Yes! :)

OpenStudy (anonymous):

okay could you help me how to create the equation for this problem? Use the matrix method to solve the system of equations. Investor Company loaned out a total of $36,000, part at 6% interest and part at 9% interest. They reported that the annual earnings from both investments were the same amount that would have been earned by the total loan if it had been invested at 8%. Find the amount loaned at each rate. Use matrix methods. Amount loaned at 6% = $_____ Amount loaned at 9% = $_____ I`m sorry I dont get how to do this one :S @Calcmathlete

OpenStudy (anonymous):

Would you open the new post per question :)

OpenStudy (anonymous):

Done :P